OpenStudy (anonymous):

Use the distributive property to multiply 108 = (x) * (x+1) * (x+8)

OpenStudy (ghazi):

first multiply x to (x+1) then multiply the product to (x+8) ...like x*(x+1)= x*x+x*1=x^2+1

OpenStudy (ghazi):

then (x^2+x)*(x+8)= x^3+8x^2+x^2+8x= x^3+9x^2+8x

OpenStudy (anonymous):

Thanks! So if I had to put 108 = x^3 +9x^2 +8x in standard form, would it be 0 = x^3 +9x^2 +8x -108?

OpenStudy (ghazi):

yea that's right